### v2

# Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need

Abstract
We use instruction-tuned Large Language Models (LLMs) such as GPT-4, MiXtral, and Llama 3 to position political texts within policy and ideological spaces. We directly ask the LLMs where a text document or its author stand on the focal policy dimension. We illustrate and validate the approach by scaling British party manifestos on the economic, social, and immigration policy dimensions; speeches from a European Parliament debate in 10 languages on the anti- to pro-subsidy dimension; Senators of the 117th US Congress based on their tweets on the left-right ideological spectrum; and tweets published by US Representatives and Senators after the training cutoff date of GPT-4. The correlation between the position estimates obtained with the best LLMs and benchmarks based on coding by experts, crowdworkers or roll call votes exceeds .90. This training-free approach also outperforms supervised classifiers trained on large amounts of data. Using instruction-tuned LLMs to scale texts in policy and ideological spaces is fast, cost-efficient, reliable, and reproducible (in the case of open LLMs) even if the texts are short and written in different languages. We conclude with cautionary notes about the need for empirical validation.
Keywords LLM |||| ideology |||| scaling |||| text as data
1 Introduction
Much research in the social and political sciences involves estimating the positions of actors, such as politicians or political parties, in latent ideological and policy latent spaces, such as the left-right or liberal-to-conservative continuum. Widely used approaches involve the automatic processing of text documents produced by these actors, such as party manifestos (Laver et al., 2003; Slapin and Proksch, 2008) or legislative speeches (Lauderdale and Herzog, 2016). Other approaches involve human coding by experts (e.g., the Comparative Manifesto Project ) or crowd workers (Benoit et al., 2016). Yet other approaches rely on other inputs, such as roll call votes (Poole and Rosenthal, 1985), Twitter connections (Barberá, 2015), or campaign donations (Bonica, 2014).
In this article, we propose a new approach to locate text documents in ideological and policy spaces using instruction-tuned Large Language Models (LLMs) and evaluate its performance. These models are LLMs optimized for dialogue use cases and are typically interacted with via chatbots such as ChatGPT. We focus on four scaling tasks using texts of different types, contexts, and lengths. First, we position political texts on the economic, social, and immigration policy dimensions based on party manifestos, a type of political text widely used in comparative politics research. Second, we position speeches by EU legislators, in 10 different languages, about a policy proposal on the ‘anti-subsidy’ to ‘pro-subsidy’ scale. Third, we position senators of the 117th US Congress on the left-right ideological spectrum based on the tweets they published during the Congress session. Fourth, we position individual tweets on the left-right spectrum that were published by US Representatives and Senators after the training cut-off date of the GPT-4 version used in these analyses (gpt-4-0613). [^1] The other LLMs with comparable performance analyzed in this research note were released in December 2023 or later and their training cutoff dates are not publicly available. This allows us to assess the performance of a modern LLM on prediction data we are certain were not part of the LLM pre-training data.
[…]
The scaling approach we propose is based on the direct query method introduced by Le Mens et al. (2023a), who used GPT-4 to measure the typicality of text documents in concepts. That article focused on several empirical settings, including the typicality of tweets in political parties (Democratic and Republican parties). We extend this previous work in four ways. First, we conduct a different task: we use LLMs for scaling text documents (party manifestos and legislative speeches) on particular policy dimensions rather than to measure their semantic similarity to a concept. Second, we use not only GPT-4 but also other LLMs, including open models (with architecture and weights available for download and use under permissive licenses). This overcomes the limitations regarding the future reproducibility of results produced with GPT-4 since the future availability of the version of the model used for research is not guaranteed by OpenAI. Third, we expand the text data and compare the validity of our estimates to those obtained using other scaling methods (e.g., the position of senators based on voting behavior). Fourth, we conduct performance evaluations not only on English texts but also in a multilingual context (EU legislative speeches).
We know of only two recent studies that have employed instruction-tuned LLMs (ChatGPT) to estimate policy and ideological positions. Our approach to scaling differs from both. The first study used the LLM as a probabilistic text classifier to obtain the posterior probability that a sentence in a party manifesto is “Conservative” or “Liberal” and defined the position of the manifesto as the average (across sentences) of the difference between these probabilities. Our approach directly asks the LLM for the position of text on the focal dimension, and we find that it performs better. The second study directly asked ChatGPT to compare pairs of politicians on a particular dimension and used these pairwise comparisons to construct estimates of politician position on unidimensional scales (e.g., gun control support) (Wu et al., 2023). We ask LLMs to position political texts produced by political actors instead of asking them to position political actors based on their names. Our approach is thus applicable even for political actors who are not sufficiently talked about in the data used for pre-training the LLM to reliably answer questions about them.
2 Methods and data
Table 1: LLMs used for the comparative analyses.
Table 1 lists the LLMs we used for text scaling, including their open or closed status. These consist of a set of the most recent and largest LLMs available at the end of April 2024.222We also experimented with other open models (Mistral 7B, Meta LLama 2 chat 7B, 13B, 70B, Google Gemma 1.1 7B. These performed significantly worse than the open models included here (table A1 in the appendix).
To obtain a position estimate of a text document from an LLM, we submitted a prompt that contained a ‘user message’ that instructed it to return such an estimate. For example, to locate an excerpt of a party manifesto on the pro- to anti-immigration scale, we used:
You will be provided with a text from a party manifesto. $\ll$ Text from the party manifesto $\gg$ Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
In all cases, we set the temperature parameter to 0, to ensure that the LLM would generate its response by selecting the most likely next token, and thus make the LLM responses as deterministic as possible (to ensure replicability). We also set the maximum number of tokens in the response to 20. This parameter does not affect the nature of the message returned by LLMs; it cuts the response down to 20 tokens if the LLM intended to generate a longer response. This ensures speed (token generation tends to be relatively slow) and limit costs (pay-per-use APIs charge per token submitted in the prompt and per token returned in the response). Finally, whenever this option was available, we set the response format to be a JSON object.
When the combination of text document and instructions resulted in a prompt longer than 4,096 tokens, we split the text document in chunks such that the prompts submitted to LLMs would not be longer than 4,096 tokens. The models we used can theoretically accept much longer prompts (up to 128K tokens for GPT-4 Turbo), but our tests revealed that submitting longer prompts leads to poorer performance. We obtained the position of each part of the text and took the weighted average of the scores for each part as the position of the text document (where the weight of a part was proportional to its number of tokens).
The appendix and the code shared on the companion OSF project provide the exact prompts and further details.
2.1 Data
2.1.1 British party manifestos and EU policy speeches in 10 languages
We obtained the texts, expert survey position estimates, and crowd-sourced position estimates from the replication package of Benoit et al. (2016). We scaled the British party manifestos analyzed in that paper on the same policy dimensions (economic and social policy for 18 manifestos, immigration policy for 8 manifestos). We also scaled (on the pro- to anti-subsidy dimension) the 36 speeches of a European Parliament debate on state subsidies. These were delivered in 10 different languages.
2.1.2 Senators of the 117th Congress
We obtained the list of senators from VoteView.com, their Twitter usernames, and downloaded the tweets they published during the Congress session through the Twitter API. Those who published fewer than 100 tweets during the Congress session were excluded from the analysis, resulting in a set of 98 senators.
2.1.3 Tweets published by US Congress members after the training cut-off date of GPT-4
We used the 900 tweets originally used in Le Mens et al. (2023a). In November 2023, we recruited 597 Prolific participants to each rate 30 tweets by answering the following question: “Where does this text stand on the ‘left’ to ‘right’ wing scale? If the text does not have political content, select ‘Not Applicable’.” Participants were not given any instructions as to what we mean by ‘left’ or ‘right.’ The crowdsourced position estimate of a tweet is the average of these ratings. All tweets, except one, received at least one position rating (different from ‘NA’), leading to a test data set of 899 tweets and their crowdsourced position estimates.
3 Results
3.1 British party manifestos
Figure 1: Positioning British party manifestos on specific policy dimensions: Pearson correlations between the benchmark (expert survey placement analyzed in Benoit et al. (2016)) and either position estimates obtained with LLMs or fine-tuned BERT, or crowdsourced position estimates reported in Benoit et al. (2016). The numbers in parentheses indicate the number of documents for which the focal LLM returned a numeric position score.
The position estimates obtained with the highest performing LLMs are very highly correlated with the Expert Survey Placements, more so than the crowdsourced position estimates (Figure 1). These results were obtained without providing any description of the policy dimension to the LLMs. Similar results were obtained by including such descriptions (Figure A5).
To put these results into perspective, we used the crowdworkers’ position ratings collected by Benoit et al. (2016) to train a supervised probabilistic text classifier based on BERT, a transformer language model from an earlier generation than LLMs analyzed here (Devlin et al., 2018). We used the trained classifier to obtain position estimates of the manifestos’ sentences and, in turn, of the manifestos. This approach did not yield better results than those obtained with the best LLMs, although the latter were (most likely) not specifically trained to scale these party manifestos.
3.2 EU legislative speeches in 10 Languages
We obtained position estimates of each speech on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension, with instructions including background information.We used the original text of each speech, without translation or any other processing. Benoit et al. (2016) obtained 6 crowdsourced position estimates for each speech from crowdworkers coding the official translations in English, German, Greek, Italian, Polish, and Spanish. For each speech, we took the simple average of the six position estimates as a benchmark (Cronbach’s alpha reliability of .99). The correlation between this benchmark and the position estimates obtained with LLMs on the speeches in their original languages is .89 or higher for the highest performing LLMs (see upper right panel of Figure 2).
Refer to caption              Refer to caption Refer to caption
Figure 2: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension: Pearson correlations between the benchmark (crowdsourced position estimates reported in Benoit et al. (2016)). The numbers in parentheses indicate the number of documents for which a position was obtained.
For the best LLMs, the performance remains essentially the same when using official translations for both LLM and crowdsourcing, demonstrating that these LLMs can be used for scaling untranslated texts in multilingual contexts. As a performance benchmark, we report the pairwise correlations between crowdsourced estimates obtained from the translations of speeches in the different languages (upper left panel of Figure 2). These correlations are of similar magnitude.
3.3 Senators of the 117th US Congress
Figure 3: Positioning Senators of the 117th Congress on the left-right ideological spectrum based on their tweets ($N=98$). Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others).
We queried the LLMs to obtain the positions, on the left-right ideological spectrum, of random samples of 100 tweets published by each senator during the 117th Congress session (January 3, 2021 to January 3, 2023). The position estimate of a senator is the average position of their tweets. Figure 3 shows that the resulting position estimates are highly correlated with those based on the roll call votes of the senators during that Congress session (Nokken and Poole, 2004), overall, and within-party. These position estimates are also highly correlated with those based on campaign funding (2020 CF scores, Bonica (2014)), but less so within party (Figure A14). Scaling the aggregated text of the 100 tweets in one or two prompts of up to 4,096 tokens produced similar results (Figure A15).
To compare these results with those obtained through approaches that do not require the submission of prompts to an LLM, we constructed position estimates for each tweet by computing their typicality in the Republican Party and their typicality in the Democratic Party by training probabilistic text classifiers (Le Mens et al., 2023b). We then defined the position of a tweet as the difference between these two typicalities. We used text classifiers based on fine-tuned BERT (the highest-performing approach in Le Mens et al. (2023b)), fine-tuned GloVe word embeddings, and a naive Bayes classifier based on word frequencies (TF-IDF) trained on 220K tweets by senators of the 116th and 117th US congresses. None of these approaches matches the best-performing LLMs, especially when it comes to capturing within-party differences.
3.4 Tweets published by members of the US Congress after the training cut-off date of GPT-4
Figure 4: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$).
The LLMs did not return position estimates for all tweets but returned ‘NA’ for a subset of tweets, indicating that they judged that these tweets did not have enough political content to return a position estimate (see Appendix H for further discussion). The correlations between the position estimates produced by the best LLMs and crowdsourcing are very high, as shown in Figure 4. Position estimates reflect differences between-party and within-party. They are also higher than those produced by supervised classifiers trained on the approximately 1 million tweets published by members of the US Congress during the 116th and 117th Congress sessions.333 See  Appendix H.2 for an approach that asks LLMs to return typicality ratings. These results are all the more impressive considering that the comparison (crowdsourced position estimates) involves a benchmark of imperfect reliability (split-half correlation with Spearman-Brown correction; Overall: .95; Tweets by Democrats: .83; Tweets by Republicans: .86).
4 Discussion
These results demonstrate that “just asking” a modern instruction-tuned LLM for the position of text documents in ideological and policy spaces can produce valid position estimates, even with very short text documents such as tweets.
This approach has the potential to greatly expand the scope of text analysis due to its accuracy, low cost, speed, ease of implementation and reproducibility (for open LLMs). Querying instruction-tuned LLMs is much less costly than human coding even when using pay-per-use APIs. Using GPT-4 to scale the 900 tweets analyzed in section 3.4 cost \$1.5 versus £1,626 for the crowdsourced position estimates; and running Meta Llama 3 on a laptop is free (though not as fast). Moreover, the accuracy of the position estimates produced with LLMs compares well to those produced by human coders. With the party manifestos, the position estimates of the best LLMs perform better than the crowdsourced position estimates, for all three policy dimensions. And with individual tweets, analyses reported in Appendix I show that the position estimates returned by the best LLMs are as accurate as the average of the ratings of about 4 or more independent human coders.
We have shown that some LLMs can reliably scale texts in their original languages, making the proposed approach particularly valuable for comparative research. But some LLMs with multilingual capabilities perform better in English than in other languages, probably because their pre-training data mainly consists of English-language texts. [^4] See for example, https://mistral.ai/news/mixtral-8x22b/. Systematic differences in accuracy across languages could lead to systematically different results in downstream analyses (e.g., a regression could lead to significant results in the English language setting, but not in the French language setting) simply due to the difference in the precision of the estimates obtained in the different languages. Addressing issues of differential measurement errors, bias, and their implications for downstream inference is necessary for LLMs to realize their full potential in comparative research.
Which LLM should researchers who want to “ask an LLM” for the position of text documents in ideological and policy spaces choose? Five main considerations come into play: performance with respect to a relevant benchmark, cost, speed, data protection, and reproducibility. Until recently (Autumn 2023), commercial LLMs such as GPT-4 had a significant performance advantage on scaling tasks compared to the best open LLMs (Mistral 7B, Llama-2), but did not guarantee future reproducibility and involved sending data to OpenAI’s servers, creating potential data protection issues. This performance gap with the best open models has essentially vanished as of April 2024, at least in the comparative analyses we have performed for this research note. Therefore, we recommend that researchers use open LLMs such as Mixtral (8X22B) and Meta Llama 3 (70B) for text scaling tasks. They are cheaper and, most importantly, their architecture and weights are freely available for download, which guarantees a high level of data protection (if executed locally on a computer with enough memory) and the reproducibility of research results.
“Asking an LLM might be all you need” to obtain accurate scaling results, but some caution is in order. Empirical validation and critical reflection on the scope of application of validation results are necessary steps for a scientifically valid use of LLMs for scaling and other measurement tasks in social and political science (for best practices, see the guidelines by Törnberg (2024)). The excellent performance of LLMs reported in this research note could tempt readers to use this approach in other domains while skipping the validation stage. But we advise them against doing so. LLMs are well-known for generating biased and unreliable results in some empirical settings. Until other researchers have shown that “just asking” a particular LLM provides accurate scaling results in a variety of empirical settings and focal latent dimensions, we cannot be certain about the breadth of settings in which LLMs perform well for scaling tasks and, a fortiori, other measurement or coding tasks. Moreover, evidence that an LLM performs well does not imply that others of the same generation will also perform well. We replicated some of the above analyses with Google Gemma 1.1. 7B (released in April 2024) and obtained much worse performance than with other recent LLMs. Until more is known about the range of domains in which LLMs perform well at scaling and other measurement tasks, case-by-case empirical validation remains essential.
5 Data Availability
Data and scripts have been deposited in OSF (https://osf.io/x2u5m/).
Appendix
Prepared for "Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need" by Gaël Le Mens and Aina Gallego
Appendix Appendix A Data
Appendix A.1 British party manifestos
We downloaded the replication materials for Benoit et al. (2016) from https://github.com/kbenoit/CSTA-APSR/tree/master (retrieved September 22, 2023). These consist of a folder CSTA-APSR-master from which we obtained the text of the party manifestos and the position estimates provided by expert surveys, sentence scaling by experts, and crowdsourced sentence scaling.
The texts of the party manifestos used for scaling on the economic and social dimension were obtained from the file CSTA-APSR-master/DataCreated/master.sentences.csv which consisted of a table of sentences. We concatenated the sentences to construct the texts of the party manifestos.
We used the same procedure for the texts of the party manifestos used for scaling on the immigration dimension from the file CSTA-APSR-master/DataCF jobs/immigration/
CFjobsresultsImmigration/f354277_immigration1.csv.
Appendix A.2 EU legislative speeches
We used the text of the speeches available in the replication package for Benoit et al. (2016) (in folder 07 Additional Texts). To construct the crowdsourced position estimates of the speech based on translations in English, German, Greek, Italian, Polish, and Spanish, we used the raw crowdsourcing coding files (fEnglishCombined.csv, f655316.csv, f656054.csv, f656640.csv, f658174.csv, f658191.csv).
Appendix A.3 Senators of the 117th US Congress
We obtained the list of senators from VoteView.com and the Twitter usernames of most senators from the data published with Le Mens et al. (2023a) (lattest_legislators_info.csv). We complemented the list of usernames with a manual search on Twitter. We deposited a table that provides the Twitter usernames we used for each of the senators in the replication folder. (Twitter developer terms prevent us from posting the Tweets we used to position the senators but allow us to post the Tweet IDs of the tweets so that users with access to the API can easily re-download them).
We downloaded the tweets they published during the Congress session via the Twitter API (in 05/2023 and 11/2023).
For each senator, we randomly sampled 100 tweets (with publication dates between January 3 and January 3, 2023). Those who published fewer than 100 tweets during the congressional session were excluded from the analysis, resulting in a set of 97 senators.
Appendix A.4 Tweets published by members of the US Congress after the training cut-off date of GPT-4 – obtaining crowdsourced position estimates
In November 2023, 597 Prolific participants (US residents, average age 40 years, 52% male, 46% female, 2% others) each rated 30 tweets. We used the same 30 sets of 30 tweets, composed of 15 tweets from each party as in Le Mens et al. (2023a). Participants were randomly assigned to one of these sets. We obtained between 19 and 21 position ratings for each tweet, with an average of 20 ratings per tweet. The crowdsourced position estimate of a tweet is the average of these ratings. All tweets, except one, received at least one position rating (different from ‘NA’), leading to a test data set of 899 tweets and their crowdsourced position estimates.
To provide their position ratings, participants responded to the question: “Where does this text stand on the ‘left’ to ‘right’ wing scale? If the text does not have political content, select ‘Not Applicable’.” Participants were not given any instructions as to what we mean by ‘left’ or ‘right.’ They reported their scores using a slider with endings ‘Extremely Left’ (left side, coded 0) to ‘Extremely right’ (right side, coded 100). The slider was centered when the page appeared on the screen. Participants could use a checkbox to respond ‘Not Applicable’.
Importantly, participants were not provided with any information about the author of the tweet — just the text of the tweet. Therefore, they were not informed whether the tweet was written by a Democrat or a Republican.
To ensure response quality, we screened Prolific participants based on the following criteria:
  • Fluent languages include English;
  • Approval Rate: Minimum 95%;
  • Country of Residence: US;
  • Political Spectrum (US): they should have provided a response to this question (the response could be anyone of {Conservative, Moderate, Liberal, Other}),
  • U.S. Political Affiliation: they should have provided a response to this question (the response could be anyone of {Democrat, Republican, Independent, Other, None}).
The last two items were included to increase the likelihood that participants would have some knowledge of US politics. After providing informed consent, participants were asked to reflect on left- and right-wing tweets. They answered one question about each side of the ideological spectrum:
What do you expect to see in a tweet that expresses a ‘left’ [‘right’] wing ideology? You could write about possible topics addressed by such tweets, and or opinions you expect the authors to have about these topics (min. response length: 100 characters).
The display order of the two questions was randomized among the participants. The responses to these questions were not analyzed (the purpose of these questions was to make participants think about the left-to-right ideological spectrum before they would respond to the scaling questions.
On the next page, they received short instructions about the position rating task.
For each of the 30 tweets, you will be asked the following question: Where does this tweet stand on the ‘left’ to ‘right’ wing scale? You will report your response using a continuous slider that goes from ‘Extremely left’ to ‘Extremely right’. There is no right or wrong answer. We are interested in your subjective opinion. If the text does not have political content, select ‘Not Applicable’. Note that the slider will appear on the screen 6 seconds after the text of the tweet.
Then they looped through the 30 tweets and provided their position ratings for each of them.
Finally, the study ended with a short demographic questionnaire and some questions about their political orientation, their political opinions, and their frequency of use of Twitter (not analyzed in this article).
Refer to captionRefer to caption
Figure A1: Tweets published by members of the 118th US Congress: Left panel:Histograms of responses by human respondents to the question about the position of each tweet on the Left-Right ideological spectrum (left panel). Right panel: Histogram of crowdsourced position estimates.
Appendix Appendix B Constructing and submitting prompts to LLMs
We used different functions to interact with the different models (see below). Most of these functions use the same parameters and parameter values:
  • messages: a Python list of dictionaries (with just one 1 element in our case) that contains the prompt (see details below).
  • max_tokens: 20. This parameter cuts the response of the LLM to a maximum of 20 tokens. It does not affect the nature of the response (it is not an instruction to respond in a few tokens). We used this small value to limit cost (pay-per-use APIs charge per number of tokens in the prompt and the response) and favor speed (the token generation process is relatively slow).
  • temperature: 0. This implies that when the model generates tokens in response to the prompt, it picks the most likely token, given the posterior distribution of candidate tokens. This ensures that the output is as deterministic as possible (there can be some randomness if there are ties).
  • top_p: 1. This parameter is essentially irrelevant when temperature is set to 0.
  • n: 1. This parameter sets the number of responses returned by the OpenAI API in response to the prompt. (Other models do not have this parameter).
In all cases, messages is defined as follows:
messages=[{
    "role": "user",
[…]
ΨΨΨ]
where text is the text document to be scaled (e.g., tweet, set of tweets, part of a party manifesto or speech) and system_message is a Python dictionary with two entries:
  • pre_text: Text to be included in the prompt before the text to be scaled.
  • post_text: Text to be included in the prompt after the text to be scaled.
Therefore, content is a string variable that contains the text to be scaled and the scaling instructions submitted to the LLM. In some cases, it is further processed by a utility function to add some special tokens necessary for a particular LLM.
Appendix B.1 OpenAI API
  • GPT-4 (gpt-4-0613): We used the following command:
    response_raw = client.chat.completions.create(
    ΨΨΨΨΨ  model=llm,
    ΨΨΨΨΨ  messages = messages,
    ΨΨΨΨΨ  max_tokens= max_tokens,
    ΨΨΨΨΨ  temperature=temperature,
        ΨΨΨΨ  top_p = top_p,
    ΨΨΨΨΨ  n = n
    ΨΨΨΨΨ)
  • GPT-4 Turbo (gpt-4-0125-preview) and GPT 3.5 Turbo (gpt-3.5-turbo-0125): We used the following command (a JSON mode is available for these models but not for the earlier models):
    response_raw = client.chat.completions.create(
        model=llm,
        response_format={ "type": "json_object" },
        messages=messages,
        max_tokens= max_tokens,
        temperature=temperature,
        top_p = top_p,
        n = n
Appendix B.2 Mistral AI API
The variable messages is processed by a useful function part of the mistralai Python package before submission to the LLM via the API.
messages_Mistral_API = [ChatMessage(role="user", content=messages[0][‘content’])]
  • Models published in 2023: open-mistral-7b,open-mixtral-8x7b: We used the following command from the mistralai Python package:
    response_raw = client.chat(
        model=llm,
        messages=messages_Mistral_API,
        max_tokens = max_tokens,
        temperature = temperature,
        top_p = top_p,
  • Models published in 2024: mistral-small-2402, mistral-large-2402, open-mixtral-8x22b: We used the following command from the mistralai Python package (a JSON mode is available for those model but not for the models published in 2023):
    response_raw = client.chat(
        model=llm,
        response_format={ "type": "json_object" },
        messages = messages_Mistral_API,
        max_tokens = max_tokens,
        temperature=temperature,
        top_p = top_p,
Appendix B.3 Running models with the llama-cpp-python Python library
Some LLMs are available for download in .gguf format, in particular quantized models (with compressed weights) that can be run locally on the user’s computer (with Apple Silicon or a powerful consumer-grade GPU such as Nvidia RTX3090). These models are executed with the llama-cpp-python Python library. Prompts need to be formatted for each model (by adding some special tokens before and after the main message, and that differ across models). To construct the prompt from the message, we used Huggingface’s transformer’s library. The Huggingface website contains a vast catalogue of LLMs, with model architecture, weights, and associated utilities. By pointing the transformer library to the URL of a particular LLM (hf_path below), the user can access utility functions adapted to that LLM (tokenizer.apply_chat_template). We used the following commands:
tokenizer = AutoTokenizer.from_pretrained(hf_path)
prompt = tokenizer.apply_chat_template(messages,
    tokenize=False,
    add_generation_prompt=True
model_local = Llama(model_path = model_path,
    n_ctx = config_LLM.max_tokens_text_scaling,
    n_gpu_layers = -1,
    use_mlock = True,
    logits_all = False,
response_raw = model_local(
    prompt = prompt,
    max_tokens = max_tokens,
    temperature=temperature,
    top_p=top_p,
Here Llama is a function that loads a model stored locally (at path model_path into the function model_local. Then the prompt is submitted to the LLM via this function.
Appendix B.4 Running models on an Apple computer with the mlx-lm Python library
The mlx-lm Python library is a deep learning framework optimized for Apple computers with ‘Apple Silicon’ processors (e.g. variants of M1, M2, M3). It plays the same role for these processors as that of the widely used Pytorch library for the NVidia GPUs.
The LLMs executed with the mlx-lm Python library were downloaded from HuggingFace.co. With each model comes a tokenizer that, among other things, can convert a prompt to the format adapted to the particular model used for scaling. The prompt query was then submitted to the LLM via the generate function of the mlx-lm library.
model_mlx, tokenizer = load(mlx_path)
prompt = tokenizer.apply_chat_template(
    messages, tokenize=False, add_generation_prompt=True
response_raw = generate(model_mlx,
    tokenizer,
    prompt=prompt,
    max_tokens = max_tokens,
    temp = temperature,
    top_p=top_p,
    verbose=True)Ψ
Here the model is stored locally in folder mlx_path. The function load from the mlx-lm library loads the LLM in the model_mlx variable and the tokenizer in the tokenizer variable. Then the tokenizer is used to pre-process the messages to create the prompt. And finally, the prompt is submitted to the LLM via the generate function from the mlx-lm library.
Appendix Appendix C Messages submitted to LLMs
In all cases, messages is defined as follows:
messages=[{
    "role": "user",
[…]
ΨΨΨ]
Next, we report on the values of pre_text and post_text in the various settings.
Appendix C.1 British party manifestos - scaling without explanations about the policy dimensions.
Appendix C.1.1 pre_text
You will be provided with a text from a party manifesto.
Appendix C.1.2 post_text
  • Economic policy:
    Where does this text stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. To compute the score, start by first identifying the parts of the text that are about economic policy. Then, compute the score based on these parts of text. If the text does not refer to economic policy, return "NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
  • Social Policy:
    Where does this text stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. To compute the score, start by first identifying the parts of the text that are about social policy. Then, compute the score based on these parts of text. If the text does not refer to social policy, return "NA"". You will only respond with a JSON object with the key Score. Do not provide explanations
    Immigration policy:
    Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.2 British party manifestos - scaling with explanations about the policy dimensions.
Appendix C.2.1 pre_text
You will be provided with a text from a party manifesto.
Appendix C.2.2 post_text
We used post_text that included explanations about the scale dimension and the two ends of the scale designed to be as close as possible to the instructions given to crowdworkers in Benoit et al. (2016).
  • Economic policy:
    Where does this text stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? “Economic” policies deal with all aspects of the economy, including: Taxation; Government spending; Services provided by the government or other public bodies; Pensions, unemployment and welfare benefits, and other state benefits; Property, investment and share ownership, public or private; Interest rates and exchange rates; Regulation of economic activity, public or private; Relations between employers, workers and trade unions. “Left” economic policies tend to favor one or more of the following: High levels of services provided by the government and state benefits, even if this implies high levels of taxation; Public investment. Public ownership or control of sections of business and industry; Public regulation of private business and economic activity; Support for workers/trade unions relative to employers. “Right” economic policies tend to favor one or more of the following: Low levels of taxation, even if this implies low levels of levels of services provided by the government and state benefits; Private investment. Minimal public ownership or control of business and industry; Minimal public regulation of private business and economic activity; Support for employers relative to trade unions/workers. Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. To compute the score, start by first identifying the parts of the text that are about economic policy. Then, compute the score based on these parts of text. If the text does not refer to economic policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
  • Social Policy:
    Where does this text stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? “Social” policies deal with aspects of social and moral life, relationships between social groups, and matters of national and social identity, including: Policing, crime, punishment and rehabilitation of offenders; Immigration, relations between social groups, discrimination and multiculturalism; The role of the state in regulating the social and moral behavior of individuals. “Liberal” social policies tend to favor one or more of the following: Recording the contents of political text on economic and social scales; Policies emphasizing prevention of crime, rehabilitation of convicted criminals; The right of individuals to make personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies penalizing discrimination against particular social groups and/or favoring a multicultural society. “Conservative” social policies tend to favor one or more of the following: Policies emphasizing more aggressive policing, increasing police numbers, conviction and punishment of criminals, building more prisons; The right of society to regulate personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies favoring restriction of immigration, and/or opposing explicit provision of state services for minority cultures. Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. To compute the score, start by first identifying the parts of the text that are about social policy. Then, compute the score based on these parts of text. If the text does not refer to social policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations
    Immigration policy:
    Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? “Immigration” policy relates to all government policies, laws, regulations, and practices that deal with the free travel of foreign persons across the country’s borders, especially those that intend to live, work, or seek legal protection (asylum) in that country. Examples of specific policies that pertain to immigration include the regulation of: work permits for foreign nationals; residency permits for foreign nationals; asylum seekers and their treatment; requirements for acquiring citizenship; illegal immigrants and migrant workers (and their families) living or working illegally in the country. It also includes favorable or unfavorable general statements about immigrants or immigration policy, such as statements indicating that immigration has been good for a country, or that immigrants have forced local people out of jobs, etc. Pro-immigration policies Examples of “pro” immigration positions include: Positive statements about the benefits of immigration, such as economic or cultural benefits; Statements about the moral obligation to welcome asylum seekers; Policies that would improve conditions for asylum seekers and their families; Urging an increase the number of work permits for foreign nationals; Making it possible for illegal immigrants to obtain a legal status or even citizenship; Reducing barriers to immigration generally. Anti-immigration policies Examples of “anti” immigration positions include: Negative statements about consequences of immigration, such as job losses, increased crime, or destruction of national culture; Arguments about asylum seekers abusing the system; Policies to deport asylum seekers and their families; Urging restrictions on the number of work permits for foreign nationals, including points systems; Deporting illegal immigrants and their families; Increasing barriers to immigration generally. Neutral immigration policy statements Examples of neutral statements about immigration policy: Advocating a balanced approach to the problem; Statements about administrative capacity for handling immigration or asylum seekers; Statements that do not take a pro- or anti-immigration stance generally, despite making some statement about immigration. Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return “NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.3 EU legislative speeches - scaling with background explanations.
Appendix C.3.1 pre_text
You will be provided with a text from a debate over policy in the European parliament.
Appendix C.3.2 post_text
We used the following post_text that included background information about the context of the legislative debate (especially the nature of the proposal being discussed), the scale dimension and the two ends of the scale designed to be as close as possible to the instructions given to crowdworkers in Benoit et al. (2016) (note that we reverse-coded the results in the analyses, to be consistent with the scale direction in Benoit et al. (2016)):
Where does this text stand on the ‘pro-subsidy’ to ‘anti-subsidy’ scale, in terms of state subsidies for uncompetitive coal mines? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-subsidy’ and 100 means ‘Extremely anti-subsidy’ ## Background. The debates are taken from a debate in the European Parliament over the ending of state support for uncompetitive coal mines. In general, state aid for national industry in the European Union is not allowed, but exceptions are made for some sectors such as agriculture and energy. At stake here were not only important policy issues as to whether state intervention is preferable to the free market, but also the specific issue for some regions (e.g. Ruhrgebiet in Germany, the north-west of Spain, the Jiu Valley in Romania) where the social and economic impacts of closure would be significant, possibly putting up 100,000 jobs at risk when related industries are considered. Specifically, the debate concerned a proposal by the European Commission to phase out all state support by 2014. Legislation passed in 2002 that allowed for state subsidies to keep non- profitable coal mines running was due to end in 2010. The Commission proposed to let the subsidies end, but to allow limited state support until 2014 in order to soften the effects of the phase-out. A counter proposal was introduced to extend this support until 2018, although many speakers took the opportunity to express very general positions on the issue of state subsidies and energy policy. ## The ‘pro-subsidy’ to ‘anti-subsidy’ scale. The “anti-subsidy" position was the essence of the Commission proposal, which would have let subsidies end while offering limited state aid until 2014 only. Examples of “anti-subsidy" positions: Statements declaring support for the commission position. Statements against state aid generally, for reasons that they distort the market. Arguments in favor of the Commission phase-out date of 2014, rather than 2018. In the strong form, the “pro-subsidy" position, involved rejecting the Commission proposal and favoring continuing subsidies indefinitely. In a weaker form, this involved supporting the compromise to extend limited state support until 2018. Examples of “pro-subsidy" positions: Arguments that keeping the coal mines open to provides energy security. Arguments that coal mines should be kept open to provide employment and other local economic benefits. Preferences for European coal over imported coal, for environmental or safety reasons. ## Instructions. To compute the score, start by first identifying the parts of the text that are about subsidy policy. Then, compute the score based on these parts of text. If the text does not refer to subsidy policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.4 EU legislative speeches - scaling without background explanations.
Appendix C.4.1 pre_text
You will be provided with a text from a debate over policy in the European parliament.
Appendix C.4.2 post_text
Where does this text stand on the ‘pro-subsidy’ to ‘anti-subsidy’ scale, in terms of state subsidies for uncompetitive coal mines? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-subsidy’ and 100 means ‘Extremely anti-subsidy’. To compute the score, start by first identifying the parts of the text that are about subsidy policy. Then, compute the score based on these parts of text. If the text does not refer to subsidy policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.5 Tweets
Appendix C.5.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ’right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set score to ‘NA’. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.5.2 post_text
Left empty.
Appendix C.6 Tweets published by each senator
Appendix C.6.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ’Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to“NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.6.2 post_text
Left empty.
Appendix C.7 Sets of tweets published by each senators
Appendix C.7.1 pre_text
You will be provided with the text of a set of tweets published by a member of the US Congress. Tweets are separated by a line break followed by the string of characters <TWEET>.
Appendix C.7.2 post_text
Where does the author of these tweets stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix Appendix D Other Scaling Methods
Appendix D.1 British party manifestos - fine-tuned BERT classifier
Note: We used a Python Jupyter notebook executed on Google Colab adapted from that made available in the companion OSF folder to the paper on using machine learning to measure typicality by Le Mens et al. (2023b) (https://osf.io/ta273/).
We implemented a BERT classifier that would predict the sentence coding crowdworkers in Benoit et al. (2016). For the manifestos used for scaling on the Economic and Social policy dimensions, we specified an 11 class sentence classifier in which the classes were the 5 positions on the Economic policy dimension [-2,-2,0,1,2] the 5 positions on the Social policy dimension [-2,-2,0,1,2], and ‘nan’ in case a crowdworker judged a sentence to be neither about economic nor social policy. We divided the data into training, validation and prediction sets, ensuring that all ratings for a given sentence were in just one of these three sets. We used 45% of the sentences for training, 5% for validation, and 50% for prediction. This produced a training set with 98K human ratings, a validation set with 10K human ratings, and a test set with 107K human ratings.
We implemented our BERT model using the bert-base-uncased, the TensorFlow machine-learning library and its higher-level wrapper Keras. We used the Adam optimizer to minimize the cross-entropy loss function to fine-tuning the model. This amounts to trying to find the parameters that maximize the likelihood of the true categories in the training data. We used Google Colab with a TPU-equipped virtual machine to train the model. We used the following (standard) parameter values:
  • Batch size: 64
  • Max number of tokens: 512 (this is the maximum possible with bert-base-uncased)
  • Optimizer: adam
  • Loss function: categorical_crossentropy
  • Learning rate: 2e-5
  • Epochs: 200 (we used a training routine that would treat this parameter as the maximum possible number of epochs - it automatically stopped the training when the validation loss started to increase)
We applied the fine-tuned model on the sentences in the prediction set. This gave a vector of categorization probabilities in the 11 candidate classes. To obtain the policy area of the focal sentence, we first added the categorization probabilities for the 5 possible dimensions of the economic policy dimension, and did the same for the Social Policy dimension. This gives a vector of 3 probabilities for Economic Policy, Social Policy and ‘Other’. We then assigned the focal sentence to the most likely policy area (including ‘Other’).
To obtain the position on the economic policy dimension, of a sentence about Economic policy (as per the result of the classification in policy areas described above), we took the weighted average of the 5 possible positions on that dimension, where the weights were the categorization probabilities in these 5 classes (rescaled to sum to 1). From this, we constructed a position estimate for each manifesto as the average position of its sentences.
We followed the same procedure to obtain position estimates of sentences and then manifestos on the Social policy dimension.
The manifestos used for scaling on the Immigration policy dimension were a separate set of party manifestos. For these, the crowdworkers coded each sentence on a 3 point dimension. We adapted the method just described (in the context of Economic and Social policy dimensions) to this context. This dataset of scaled sentences was much smaller that that for Economic and Social policy, giving a training set of about 11K observation, of validation set of about 1.2K observations and a prediction set of about 12K observations.
Appendix D.2 Tweets - fine-tuned BERT classifier
Note that for tweets, we did not train the models on data produced by human coders, but instead trained the classifiers to predict the party of the Congress member who published the tweet (Democratic Party or Republican Party), following the approach proposed by Le Mens et al. (2023b) and implemented on these Twitter data in Le Mens et al. (2023a). We used the approach used in this latter paper by slightly adapting the jupyter python notebooks made available at on the companion OSF folder (https://osf.io/ta273/).
For the tweets published by the senators of the 117th US congress, we used as a training set the tweets published by senators of the 116th and 117th Congress sessions, ensuring that none of the tweets in the prediction set (100 tweets per senator) was in the training or validation set. We used the categorization probabilities in the two parties to compute the typicality of each tweet in each party (following Le Mens et al. (2023b)) and took the difference in typicalities in the Republican and Democratic parties as the position of the tweet on the left-right ideological spectrum. The position estimate of a senator is the average position of the 100 tweets in the prediction set.
For tweets published after the training cut-off date of GPT-4, we used the same training set as Le Mens et al. (2023a) and simply computed the position of each tweet following the approach explained in the previous paragraph.
Appendix D.3 Tweets - fine-tuned GloVe classifier
Note: We used a Python Jupyter notebook executed on Google Colab adapted from that made available in the companion OSF folder to the paper on using machine learning to measure typicality by Le Mens et al. (2023b) (https://osf.io/ta273/). The following text is a minor adaptation of the description of the fine-tuned GloVe classifier in that paper.
We adapted the approach described in Le Mens et al. (2023b) to the present empirical setting. We used the same model structure as for the BERT classifier, but we used GloVe word embeddings as a language representation instead of the BERT language representation. More precisely, we used a word embedding layer with pre-trained weights in our classifiers instead of the BERT language representation. We used GLOVE word embeddings (Pennington et al., 2014) to transform text documents into vectors. GLOVE is a word-embedding model, not a text-embedding model. Consequently, we needed to combine word positions in the embedding space to create a unique position for text documents. We used the average position of the words in the text as the position of the tweet in semantic space.
We selected the top 20,000 most frequent words in the training data, then we used the Glove embedding model “glove.6B.300d” to transform each of the 20,000 words into vector positions. With this embedding, we created a basic deep-learning model consisting of:
  1. 1.
    Embedding layer
  2. 2.
    Pooling 1D layer
  3. 3.
    Dense layer (with softmax activation)
We trained the model for up to 1,000 epochs with a $32e-$ 3 learning rate using a categorical cross entropy loss function.
Appendix D.4 Tweets - fine-tuned Naive Bayes classifier with tf-idf language representation
Note: We used a Python Jupyter notebook executed on Google Colab adapted from that made available in the companion OSF folder to the paper on using machine learning to measure typicality by Le Mens et al. (2023b) (https://osf.io/ta273/). The following text is a minor adaptation of the description of the naive Bayes classifier in that paper.
We also used a standard machine-learning text classifier based on a version of the Bag-of-Words representation that weighs word frequencies by diminishing the importance of words that occur in many text documents. This approach is known as ‘TF-IDF’, or ‘term frequency-inverse document frequency’, (Jones, 1972). The classifier model is the Naive Bayes classifier (Maron, 1961). This machine-learning classifier produces categorization probabilities based on word co-occurrences. It is computationally undemanding, but its representation of text documents is not sensitive to the order of words in sentences. Also, the representation of words does not depend on their semantic similarity.
To train the classifier applied on the tweets of the senators of the 117th US Congress (Section 3.3), we selected the 5,000 more frequent words from the training data (this number is constrained by the RAM of the computer on which we ran the model training, which in this case was a virtual machine on Google Colab with 50GB of RAM). We assigned an ID to each word and transformed all the book descriptions using this ID dictionary. Finally, we fit a multinomial Naive Bayes on all the book descriptions in the training data.
To train the classifier applied to the tweets published after the GPT-4 training cut-off date (Section 3.4), we selected the 2,000 more frequent words from the training data (this number is lower than in the previous paragraph because the training data is larger in this case and thus running the model training requires more RAM for the same vocabulary size).
For the construction of the dictionary, we used the sklearn package “CountVectorizer” and to fit the model we used sklearn package “MultinomialNB.”
Appendix Appendix E Additional Results - British Party Manifestos
Appendix E.1 Scatter plots of position estimates produced by LLMs and crowdworkers versus the expert survey placement
The position estimates are those used to compute the correlations reported in the body of the article (Figure 1).
Figure A2: Position of British party manifestos on the Economic policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. The numbers next to the dots are the years of the party manifestos.
Figure A3: Position of British party manifestos on the Social policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. The numbers next to the dots are the years of the party manifestos.
Figure A4: Position of British party manifestos on the Immigration policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. All the party manifestos were from 2010.
Appendix E.2 Performance of position estimates obtained by including descriptions of policy dimensions in the prompts submitted to the LLMs.
Figure A5: Replication of Figure 1 except that a description of the policy dimensions and of its extremities as similar as possible to the instructions given to crowdworkers by Benoit et al. (2016) was included in the prompts submitted to the LLMs. The numbers in bold report the Pearson correlations between the benchmark (expert survey placement) and position estimates obtained with LLMs or crowdsourced position estimates reported in Benoit et al. (2016). The numbers in parentheses indicate the number of documents for which a position estimate was obtained (position estimates are missing if the LLM did not return a position score for any of the parts of manifestos submitted to it).
Appendix Appendix F Additional Results - EU legislative speeches in 10 Languages
Appendix F.1 Scatter plots that correspond to the results reported in Figure 2
Figure A6: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning in the original languages.
Figure A7: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in English.
Figure A8: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in German.
Figure A9: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Greek.
Figure A10: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Italian.
Figure A11: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Polish.
Figure A12: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Spanish.
Appendix Appendix G Additional Results - Senators of the 117th US Congress
Visual comparison of the performance of the various scaling methods (LLMs, fine-tuned BERT classifier, fine-tuned GloVe classifier, and naive Bayes classifier with tf-idf representation).
Figure A13: Positioning Senators of the 117th Congress on the left-right ideological spectrum based on their tweets ($N=98$). Pearson correlations between the benchmark (First dimension Nokken-Poole period-specific DW-NOMINATE score) and position estimates obtained with LLMs and other scaling methods. The numbers in parentheses indicate the number of senators for which a position estimate was obtained. The data are the same as for Figure 3.
Appendix G.1 Correlation between the position estimates of the senators (using the baseline approach described in the body of the article) and a position estimate measure based on campaign funding.
Figure A14: Replication of Figure 3 using the 2020 CFscore from the Dime database (Bonica, 2014) instead of the period-specific DW-NOMINATE score. The 2020 CFscore is a period-specific estimate of ideology based on donation received.
Appendix G.2 Scaling the aggregated text of the 100 tweets in one or two prompts of up to 4,096 tokens
We used the user message to obtain the position of each senator based on the text of the 100 tweets:
You will be provided with the text of a set of tweets published by a member of the US Congress. Tweets are separated by a line break followed by the string of characters <TWEET>. <TWEET> $\ll$ Text of tweet 1 $\gg$ … <TWEET> $\ll$ Text of tweet 100 $\gg$ Where does the author of these tweets stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
For most senators, the set of 100 tweets was too long to be included in a prompt of up to 4,096 tokens. Therefore, we scaled the full set of tweets in several queries to the LLM (using the same approach as for the party manifestos) and took the weighted average of the position scores (where the weights are proportional to the number of tokens) as the position estimate of a senator.
Figure A15: Replication of Figure 3 by submitting sets of tweets in prompts with up to 4,096 tokens rather than by scaling each tweet. Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others). The Y-axis reports, for each senator, the position estimates of senators based on random samples of 100 of their tweets published during the Congress session. The X-axis reports the first dimension Nokken-Poole period-specific DW-NOMINATE score.
Appendix Appendix H Additional Results - Tweets published by members of the US Congress after the training cut-off date of GPT-4
Visual comparison of the performance of the various scaling methods (LLMs, fine-tuned BERT classifier, fine-tuned GloVe classifier, and naive Bayes classifier with tf-idf representation).
Figure A16: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$). Pearson correlations between the benchmark (crowdsourced position estimates) and position estimates obtained with LLMs. The numbers in parentheses indicate the number of documents for which a position estimate was obtained (position estimates are missing if the LLM did not return a position score for any of the parts of manifestos submitted to it). The data are the same as for Figure 4.
The LLMs failed to return position scores for some of the tweets, possibly because these did not include enough political content. To verify that this pattern of non-response is reasonable, we computed the average number of human ratings for the two sets of tweets, for each LLM (survey respondents had the option to respond ‘Not applicable’ to the question about the position of each tweet on the Left-Right ideological spectrum, if they felt the tweet did not contain enough information for them to return a position rating). Under the hypothesis that an LLM fails to return a numeric score for tweets that did not contain enough political content, but would do so for tweets than contain enough political content, we expected that the average number of human position ratings would be larger for tweets that received a position score by the LLM than for tweets that did not receive a position score. These quantities are reported in the last two columns of table A1. For all the LLMs, the pattern is consistent with the hypothesis.
Table A1: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$). The last two columns report the average number of human position ratings for tweets that received a position score, and the same quantity for those for which the LLM did not return a position score. This table includes the LLMs from the body of the paper as well as a set of additional LLMs (from the previous generation). ‘Open’ LLMs can be downloaded on the user’s machine and run locally if the machine has enough video memory. ‘Local’ execution means that the model was run on our laptop after downloading the model ‘API’ execution means that the model was run for a fee on a machine managed by the provider of the model. ‘Cloud’ execution means that the model was downloaded and run via a cloud computing service (Google Colab) on a machine equipped with a TPU or a GPU. Models in .gguf format are versions of the model with compressed (‘quantized’) weights and were run locally with the llama-cpp-python package. In the names of these models,‘qX’ indicates the number of bits used to represent the model weights. We used minimal compression that would allow us to run the model on our laptop (an Apple Macbook Pro with an M1 Max processor with 64GB that can run models as large as 40GB). Meta Llama 3 8B, Llama 2 7B and Llama 2 13B were downloaded from Huggingface.co and were run locally without any compression with the mlx-lm Python package.
Appendix H.1 Distribution of position estimates produced by LLMs and other scaling methods
Figure A17: Tweets published by members of the 118th US Congress: Histograms of position estimates produced by the LLMs and other scaling methods.
Figure A18: Tweets published by members of the 118th US Congress: Histograms of crowdsourced position estimates for the subset (out of the total of 899 tweets) for which each scaling method produced a position estimate.
Appendix H.2 Scaling tweets on the left-right ideological spectrum by using LLMs to obtain their typicalities in the Democratic and Republican parties
We implemented the approach advocated in Le Mens et al. (2023a) to obtain typicality measures of tweets in political parties (and illustrated in that paper with GPT-4). This involves directly asking the LLMs for the typicality of each tweet in each party. We then construct a position estimate of each tweet on the left-right ideological spectrum as the difference between its typicality in the Republican party and its typicality in the Democratic party. Le Mens et al. (2023a) had found that the typicality measures obtained with GPT-4 had higher correspondence with human typicality ratings than those produced with other methods, such as a BERT classifier or classifiers based on text embeddings (ADA2 embeddings published by OpenAI) or text embeddings (GloVe) that were trained on vast amounts of data.
Appendix H.2.1 Messages submitted to LLMs.
We used the same message structure as that explained in  Appendix C except for a change to the ‘pre_text’ part of the user messages.
Appendix H.2.2 pre_text for typicality in Democratic Party.
You will be provided with the text of a tweet published by a member of the US Congress. How typical is this tweet of the Democratic Party? Provide your response as a score between 0 and 100 where 0 means ‘Not typical at all’ and 100 means ‘Extremely typical’. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix H.2.3 pre_text for typicality in Republican Party.
You will be provided with the text of a tweet published by a member of the US Congress. How typical is this tweet of the Republican Party? Provide your response as a score between 0 and 100 where 0 means ‘Not typical at all’ and 100 means ‘Extremely typical’. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix H.2.4 post_text.
Left empty.
Appendix H.2.5 Results.
The correlations with crowdsourced position estimates are similar to those obtained with the baseline approach reported in the body of the article, with some gains in terms of within-party correlation. Another key difference is that, with this approach, LLMs return position estimates for all tweets. It should be noted that there is no need for a tweet to have explicitly political content for an artificial or human judge to evaluate its typicality in a political party. This is because of a fundamental difference between this approach of constructing a position estimate and the baseline approach used in the body of the paper: the baseline approach instructs the LLM to focus on a particular dimension. The approach consists in taking the difference between typicalities in two concepts does not require the specification of a particular dimension: it amounts to a sort of projection on the direction of the vector that connects that centers of the two concepts in multidimensional semantic space.
Figure A19: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$) by asking for the typicality in the Republican party, their typicality in the Democratic party and defining the position estimate as the difference between these typicalities. The y-axes report the position of each tweet on the left-right ideological spectrum obtained with the various models. The x-axes reports crowdsourced estimates as the average position provided by participants in an online survey of US residents.
Figure A20: Tweets published by members of the 118th US Congress: Histograms of position estimates produced by the LLMs by taking the difference between the typicality in the Republican party and typicality in the Democratic party.
Appendix Appendix I Predicting the position estimates given by one human respondent to the survey of tweet positions: ENO scores
One existing characterization of the predictive performance of models of human judges focuses on predicting the ratings of one human respondent (Erev et al., 2007). It asks how many observations of other human participants are necessary to make a prediction as good as the model. The resulting number is the model’s equivalent number of observations (ENO).
Here, we predict the position ratings of one human participant based on the average of $N$ position ratings of other participants. We compute the Pearson correlation between these two quantities. The resulting correlation will depend on the specific ratings selected as criterion and predictor variables. So, for each tweet, we randomly sample one human rating as the criterion variable and treat the remaining ratings as potential predictors. We repeat this procedure 100 times and report the average correlation in Table A2. We do this for tweets that received at least 15 human position ratings (at least 15 respondents in our survey did not select ‘NA’ for that tweet), which leads to a set of 598 tweets. We repeat the procedure for values of $N$ that range from 1 to 14 (the number of human typicality ratings minus 1), and over three sets of tweets: tweets from Democratic and Republican politicians, tweets from Democratic politicians and tweets from Republican politicians.
It should be noted that the column for N=1 reports the pairwise correlation between the position ratings given by two randomly picked human judge, computed over the tweets for which the focal model returned a position score. Because the set of tweets for which each model returned a position score differs across model, this measure of pairwise correlation differs across rows in the column. The same observation applies to the other values of N.
The ENO scores for the tweet position estimates provided by the LLMs are all higher than 1 (except for MiXtral 8X7Bq6) and go up to 5 or 6 (for GPT-4, MiXtral 8X22B and Meta Llama 3 70Bq4, over all tweets). In the latter case, this means that the position scores provided by these models are as good predictors as the average position ratings given by 5 independent human participants in our survey of Prolific participants who reside in the US.
Table A2: Tweet data: Predicting one human position rating rating with position ratings by other human coders and with the position estimates provided by the LLMs)